Need to solve by substitution method.
x+5/6y-19/6=0, 6y-21=-9x

Answer by lynnlo(4176)

x=-1,,,,,Y=5
HERE IS AN EXAMPLE FOR YOU TO STUDY STEP BY STEP
2X+3Y=5
X+Y=5
STEP(1)SOLVE ONE OF THE EQUATIONS FOR EITHER X= OR Y=
WE WILL START THE SECOND EQUATION FOR Y
X+Y=5
Y=5-X
STEP(2)SUBSTITUTE THE SOLUTION FROM STEP1,INTO SECOND EQUATION
2X+3Y=5
2X+3(5-X)=5
STEP(3) SOLVE NEW EQUATION
2X+3(5-X)=5
2X+15-3X=5
-X+15=5
=10
-X=5-15
SOLVE FOR SECOND VARIABLE
Y=5-X
Y=5-10
Y=-5
SOLUTION IS X=10,,,Y=-5
I DID AN EXAMPLE ,I HOPE THIS WILL HELP YOU UNDERSTAND THE
REST OF YOUR WORK.
X